# Not All Attention is Needed: Parameter and Computation Efficient Transfer Learning for Multi-modal Large Language Models

### v1

@@ 4.3 Attention Redundancy Evaluation
In EAS, we also adopts a -arm bandit based algorithm to automatically evaluate the redundancy of MHAs in MLLMs for downstream task, similar with Wu et al. (2023b).
Concretely, given a MLLM, we first randomly sample and tune the subnetworks with PIAs as adapters for a few epochs. Afterwards, we initialize a numerical action vector, denoted as , and then we keep training and testing the MLLM with PIAs. In each evaluation step, we skip MHA modules according to , based on which the action policy at the step is obtained by
@@
Here, is the index of skipped MHA modules. Thus, the MHA with a greater is more likely to be kept.
@@
| Method | Updated Parameters | VQA | NLVR | Flickr30K | Average | ||||
@@
| Shallow Prompt (Li & Liang, 2021a) | 0.30M | 68.51 | +28.71G | 65.69 | +26.84G | 74.20/88.60 | +28.71G | 74.25 | +28.71G |
| Deep Prompt (Jia et al., 2022a) | 1.84M | 70.78 | +6.53G | 72.64 | +5.59G | 78.84/89.40 | +6.53G | 77.92 | +6.53G |
| LoRA (Hu et al., 2022) | 0.29M | 74.00 | +0.00 | 78.82 | +0.00 | 79.86/92.60 | +0.00 | 81.32 | +0.00 |
| Adapter (Sung et al., 2022a) | 5.34M | 74.70 | +1.64G | 79.93 | +1.38G | 80.38/91.90 | +1.64G | 81.73 | +1.64G |
| Scaled PA (He et al., 2022) | 3.59M | 75.11 | +1.12G | 80.38 | +0.66G | 80.40/93.20 | +1.12G | 82.27 | +1.12G |
| DAS (Wu et al., 2023b) | 5.34M | 74.80 | -11.16G | 80.11 | -5.13G | 80.12/91.80 | -11.16G | 81.71 | -9.15G |
| EAS(Ours) | 4.07M | 74.85 | -10.65G | 80.13 | -7.38G | 80.38/92.20 | -9.12G | 81.89 | -9.05G |
| EAS(Ours) | 4.07M | 74.82 | -14.53G | 80.08 | -10.19G | 79.78/90.40 | -11.90G | 81.27 | -12.21G |
5 Experiment
5.1 Datasets and Metrics
ScienceQA (Lu et al., 2022) is a multi-modal dataset for science question answering. This dataset is divided into three subsets, i.e., train, val, and test, which have , , and examples, respectively. The questions can be categorized into Natural Science (NAT), Social Science (SOC), Language Science (LAN), Text Context (TXT), Image Context (IMG), No Context (NO), Grades 1-6 (G1-6) and Grades 7-12 (G7-12), respectively. VQA2.0 (Goyal et al., 2017a) involves images, each accompanied by at least three open-ended questions. We report the test-dev results from the online evaluation 444https://eval.ai/web/challenges/challenge-page/830/overview. NLVR (Suhr et al., 2019a) is tailored for classifying triplets consisting of two images and a question, and it has examples of human-written sentences grounded with pairs of images. Flickr30K (Plummer et al., 2017) is a cross-modal retrieval dataset, comprising images collected from Flickr, each image has five reference sentences provided by human.
5.2 Implementation Details
The base model employed in our study is LaVIN (Luo et al., 2023b), a recently introduced MLLM. Following its default settings, we use LLaMA-7B (Touvron et al., 2023) as the main network, and adopt ViT-L (Dosovitskiy et al., 2021) as the image encoder. In terms of PIA, we set the hidden dimension to 32. Besides, we employ PIAs for parameter-efficient adaptations, similar to (Luo et al., 2023b; Zhang et al., 2023). These PIAs have a hidden dimension of 8, and only one path for adaptation is used, i.e., in Eq. 4. We insert these PIAs before each attention module. For redundancy evaluation, we first train the subnetworks with randomly skipped modules for 5 epochs. Then we introduce redundancy evaluation in the next 5 epochs, where 3 subnetworks are sampled every 10 training steps. In addition, we also validate the proposed EAS on a conventional VL pre-trained model named METER (Dou et al., 2022). Its attention modules in a co-attention layer are considered as independent potential skipped components. The PIAs for skipping attention modules are configured with a hidden dimension of 144, with the remaining settings following the defaults in DAS (Wu et al., 2023b).
@@ 5.2 Implementation Details
5.3 Experimental Results
5.3.1 Quantitative Analysis
Comparison with DAS. We first compare our EAS with DAS (Wu et al., 2023b) on LaVIN and METER on four VL benchmarks, of which results are given in Tab. 1 and Tab. 2. We first observe that both DAS and EAS can achieve competitive performance, while retaining parameter and computation efficiency. Speciffically, in Tab. 1, DAS can skip Transformer layers of LaVIN for better efficiency while retraining comparable performance, i.e accuracy while efficiency. Meanwhile, EAS-7B can improve the average accuracy and inference speed by and under the same parameter scale. With a little extra parameter cost, our EAS can skip MHAs while achieving better performance than both DAS and LaVIN. Notably, compared with LaVIN and DAS, EAS-7B can improve the inference speeds by and times, respectively. In Tab.2, we compare to DAS on three classic VL benchmarks, i.e. VQA, NLVR and Flickr30K. Compared with DAS, our EAS can update fewer parameters and improve inference by with only performance drop on average. Overall, these results well validate EAS towards the target of PCETL for MLLMs.
Comparison with the state-of-the-arts. In Tab. 1 and Tab. 2, we also compare EAS with more recent advancements of MLLMs and PETL methods. We can first observe that PETL methods achieve competitive or even better performance than the fully tuned methods, while retaining a very small number of parameters to update. However, the use of adapters also obviously slows down their inference speeds. For instance, the inference of LLaMA-Adapter in Tab. 1 consumes more than seconds for each example. In contrast, LoRA does not incur any additional overhead during inference, thus only taking seconds per example. Compared with these approaches, the results of EAS are more significant in both performance and efficiency. EAS-7B improve performance of LoRA while improve inference speed up to on ScienceQA. Meanwhile, in Tab. 2, we can observe that, EAS achieves performance of Scaled PA while decreasing FLOPs. Overall, these results well validate the effectiveness of EAS towards the target of PCETL for MLLMs.
@@ 5.3.1 Quantitative Analysis
| Methods | Updated Parameters | Inference Time | Average Performance |
| LaVIN-7B | 0.06% | 3.70s | 89.41 |
| EAS-7B | 0.08% | 2.08s | 89.44 |
| EAS-7B | 0.10% | 1.92s | 89.79 |
| EAS-7B | 0.11% | 1.69s | 89.98 |
| EAS-7B | 0.13% | 1.56s | 88.42 |
@@ 5.3.2 Qualitative Analysis
Examples of Science Question Answering. In Fig. 4, we visualize the examples of EAS-7B and LaVIN-7B for both language-only (left) and image-language questions (right). From these examples, we can see that both EAS-7B and LaVIN-7B correctly answer the question. Meanwhile, they all inherit the strong language ability of LLaMA (Touvron et al., 2023), and can explain the answers fluently. Notably, for the visual question (right), EAS-7B has better and more detailed explanations for the answer, e.g., “The colony’s name comes from the fact that it was founded by people from Hampshire, a county in England”. In contrast, the response of LaVIN is logically incorrect even through its answer is right. Overall, these results confirm again the merits of EAS for the efficient adaption of MLLMs.
@@ 6 Conclusion
In this paper, we propose a novel method for parameter and computation efficient tuning of Multi-modal Large Language Models (MLLMs), named by Efficient Attention Skipping. Concretely, we first reveal that not all MHAs are necessary for the efficient adaption of MLLMs, based on which EAS adopts a granular redundancy evaluation scheme. Meanwhile, to avoid the additional computation caused by the adapter-based skip connections, EAS is also equipped with a novel Propagation-of-Information Adapter (PIA), which can not only keep parameter efficiency but also can be re-parameterized into the model without extra latency. To validate EAS, we apply it to a recent MLLM called LaVIN, and conduct extensive experiments on the multi-modal benchmark ScienceQA. The experimental results show that EAS can achieve better performance than LaVIN while speeding up inference by up to 2.18 times.
7 Brader Impact
This paper address the parameter and computation efficiency of MLLMs through efficient attention skipping (EAS), which will further advance the development of MLLMs in the machine learning community. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Influence of hidden dimension
@@ Appendix B Comparison with DAS
| LaVIN-7B (Luo et al., 2023b) | 3.8M | 32.32G | 3.70s | 89.25 | 94.94 | 85.24 | 88.51 | 87.46 | 88.08 | 90.16 | 88.07 | 89.41 |
| DAS-7B (Wu et al., 2023b) | 4.2M | 32.44G | 3.44s | 88.68 | 94.94 | 86.45 | 88.03 | 86.81 | 88.92 | 90.20 | 88.00 | 89.41 |
| DAS-7B (Wu et al., 2023b) | 4.6M | 31.67G | 3.23s | 88.99 | 94.60 | 85.09 | 87.88 | 86.51 | 88.36 | 89.72 | 88.13 | 89.15 |
| DAS-7B (Wu et al., 2023b) | 5.0M | 30.91G | 3.06s | 87.30 | 93.36 | 82.36 | 86.12 | 85.97 | 85.71 | 88.18 | 85.70 | 87.29 |
| DAS-7B (Wu et al., 2023b) | 5.4M | 30.05G | 2.94s | 86.50 | 89.31 | 84.27 | 85.63 | 83.74 | 86.55 | 87.41 | 84.90 | 86.51 |
| EAS-7B(Ours) | 4.5M | 32.82G | 2.13s | 88.81 | 95.16 | 85.36 | 87.29 | 86.32 | 88.99 | 90.35 | 87.28 | 89.25 |
| EAS-7B(Ours) | 5.1M | 32.50G | 2.08s | 88.81 | 94.83 | 86.36 | 87.68 | 86.81 | 89.06 | 90.42 | 87.67 | 89.44 |
| EAS-7B(Ours) | 5.7M | 32.05G | 1.99s | 88.99 | 94.15 | 85.36 | 87.88 | 87.46 | 88.92 | 90.68 | 86.68 | 89.35 |
| EAS-7B(Ours) | 6.4M | 31.63G | 1.92s | 90.05 | 94.04 | 85.82 | 89.25 | 87.56 | 88.78 | 90.68 | 88.20 | 89.79 |
| EAS-7B(Ours) | 7.7M | 31.10G | 1.69s | 89.52 | 95.61 | 86.36 | 88.71 | 88.25 | 88.85 | 91.26 | 87.67 | 89.98 |
Appendix C The Detailed Results
@@
Appendix D The detailed skipping results
We first visualize the accelerated MLLM searched by DAS (Wu et al., 2023b) and our EAS. From Fig. 6, we can observe that the layers skipped by both DAS and EAS tend to be deep layers. While the layers that are considered to have redundancy in EAS are deeper than those in DAS. Furthermore, we can find out that the skipped modules are concentrated behind the middle of MLLM, and gradually expand forward as the number of skipped modules increases.

### v2

@@ 4.3 Attention Redundancy Evaluation
In EAS, we adopt a -arm bandit based algorithm to automatically evaluate the redundancy of MHAs in MLLMs for downstream task.
Concretely, given a MLLM, we first randomly sample and tune the subnetworks with PIAs as adapters for a few epochs. Afterwards, we initialize a numerical action vector, denoted as , and then we keep training and testing the MLLM with PIAs. In each evaluation step, we skip MHAs according to , based on which the action policy at the step is obtained by
@@
Here, is the index of the skipped MHA modules. The MHA with a greater will be kept.
@@
5 Experiment
5.1 Datasets and Metrics
ScienceQA [36] is a multi-modal dataset for science question answering. This dataset is divided into three subsets, i.e., train, val, and test, which have , , and examples, respectively. The questions can be categorized into Natural Science (NAT), Social Science (SOC), Language Science (LAN), Text Context (TXT), Image Context (IMG), No Context (NO), Grades 1-6 (G1-6) and Grades 7-12 (G7-12), respectively. VQA2.0 [13] has images, each accompanied by at least three open-ended questions. We report the test-dev results from the online evaluation 444https://eval.ai/web/challenges/challenge-page/830/overview. NLVR2 [45] is tailored for classifying triplets consisting of two images and a question, and it has examples of human-written sentences grounded with pairs of images. Flickr30K [43] is a cross-modal retrieval dataset, comprising images collected from Flickr.
5.2 Implementation Details
The base model employed in our study is LaVIN [38], a recently introduced MLLM. Following its default settings, we use LLaMA-7B [50] as the main network, and adopt ViT-L [9] as the image encoder. In terms of PIA, we set the hidden dimension to 32. Besides, we employ PIAs for parameter-efficient adaptations, similar to [38, 54]. These PIAs have a hidden dimension of 8, and only one path for adaptation is used, i.e., in Eq. 4. We insert these PIAs before each attention module. For redundancy evaluation, we first train the subnetworks with randomly skipped modules for 5 epochs. Then we introduce redundancy evaluation in the next 5 epochs, where 3 subnetworks are sampled every 10 training steps. In addition, we also validate the proposed EAS on a conventional VL pre-trained model named METER [10]. Its attention modules in a co-attention layer are considered as independent potential skipped components. The PIAs for skipping attention modules are configured with a hidden dimension of 144, with the remaining settings following the defaults in DAS [53].
5.3 Experimental Results
5.3.1 Quantitative Analysis
Comparison with DAS. We first compare our EAS with DAS [53] on LaVIN and METER on four VL benchmarks, of which results are given in Tab. 1 and Tab. 2. We first observe that both DAS and EAS can achieve competitive performance, while retaining parameter and computation efficiency. Speciffically, in Tab. 1, DAS can skip Transformer layers of LaVIN for better efficiency while retraining comparable performance, i.e accuracy while efficiency. Meanwhile, EAS-7B2 can improve the average accuracy and inference speed by and under the same parameter scale. With a little extra parameter cost, our EAS can skip MHAs while achieving better performance than both DAS and LaVIN. Notably, compared with LaVIN and DAS, EAS-7B12 can improve the inference speeds by and times, respectively. In Tab.2, we compare to DAS on three classic VL benchmarks, i.e. VQA, NLVR and Flickr30K. Compared with DAS, our EAS can update fewer parameters and improve inference by with only performance drop on average. Overall, these results well validate EAS towards the target of PCETL for MLLMs.
| NAT | SOC | LAN | TXT | IMG | NO | G1-6 | G7-12 | ||||
| Fully tuned methods | |||||||||||
| LLaVA-7B [32] | 100.00% | 2.19s | 90.36 | 95.95 | 88.00 | 89.49 | 88.00 | 90.66 | 90.93 | 90.90 | 90.92 |
| Parameter-efficient methods | |||||||||||
| LLaMA-Adapter [54] | 0.03% | 4.35s | 84.37 | 88.30 | 84.36 | 83.72 | 80.32 | 86.90 | 85.83 | 84.05 | 85.19 |
| LaVIN-7B [38] | 0.06% | 3.70s | 89.25 | 94.94 | 85.24 | 88.51 | 87.46 | 88.08 | 90.16 | 88.07 | 89.41 |
| LaVIN-7B-LoRA [17] | 0.08% | 2.22s | 84.06 | 77.95 | 85.73 | 82.84 | 74.12 | 88.64 | 83.77 | 82.20 | 83.21 |
| Parameter and computation efficient methods | |||||||||||
| DAS-7B4 [53] | 0.07% | 3.23s | 88.99 | 94.60 | 85.09 | 87.88 | 86.51 | 88.36 | 89.72 | 88.13 | 89.15 |
| EAS-7B2 (Ours) | 0.07% | 2.13s | 89.48 | 93.93 | 87.00 | 88.32 | 87.36 | 89.76 | 90.20 | 88.99 | 89.77 |
| EAS-7B12 (Ours) | 0.11% | 1.69s | 89.52 | 95.61 | 86.36 | 88.71 | 88.25 | 88.85 | 91.26 | 87.67 | 89.98 |
| Method | Updated Parameters | VQA | NLVR2 | Flickr30K | Average | ||||
@@ 5.3.1 Quantitative Analysis
| Shallow Prompt [31] | 0.30M | 68.51 | +28.71G | 65.69 | +26.84G | 74.20/88.60 | +28.71G | 74.25 | +28.71G |
| Deep Prompt [19] | 1.84M | 70.78 | +6.53G | 72.64 | +5.59G | 78.84/89.40 | +6.53G | 77.92 | +6.53G |
| LoRA [17] | 0.29M | 74.00 | +0.00 | 78.82 | +0.00 | 79.86/92.60 | +0.00 | 81.32 | +0.00 |
| Adapter [48] | 5.34M | 74.70 | +1.64G | 79.93 | +1.38G | 80.38/91.90 | +1.64G | 81.73 | +1.64G |
| Scaled PA [15] | 3.59M | 75.11 | +1.12G | 80.38 | +0.66G | 80.40/93.20 | +1.12G | 82.27 | +1.12G |
| DAS4 [53] | 5.34M | 74.80 | -11.16G | 80.11 | -5.13G | 80.12/91.80 | -11.16G | 81.71 | -9.15G |
| EAS8(Ours) | 4.07M | 74.85 | -10.65G | 80.13 | -7.38G | 80.38/92.20 | -9.12G | 81.89 | -9.05G |
| EAS10(Ours) | 4.07M | 74.82 | -14.53G | 80.08 | -10.19G | 79.78/90.40 | -11.90G | 81.27 | -12.21G |
Comparison with the state-of-the-arts. In Tab. 1 and Tab. 2, we also compare EAS with more recent advancements of MLLMs and PETL methods. We can first observe that PETL methods achieve competitive or even better performance than the fully tuned methods, while retaining a very small number of parameters to update. However, the use of adapters also obviously slows down their inference speeds. For instance, the inference of LLaMA-Adapter in Tab. 1 consumes more than seconds for each example. In contrast, LoRA does not incur any additional overhead during inference, thus only taking seconds per example. Compared with these approaches, the results of EAS are more significant in both performance and efficiency. EAS-7B12 improve performance of LoRA while improve inference speed up to on ScienceQA. Meanwhile, in Tab. 2, we can observe that, EAS8 achieves performance of Scaled PA while decreasing FLOPs. Overall, these results well validate the effectiveness of EAS towards the target of PCETL for MLLMs.
@@
| Methods | Updated Parameters | Inference Time | Average Performance |
| LaVIN-7B | 0.06% | 3.70s | 89.41 |
| EAS-7B4 | 0.08% | 2.08s | 89.44 |
| EAS-7B8 | 0.10% | 1.92s | 89.79 |
| EAS-7B12 | 0.11% | 1.69s | 89.98 |
| EAS-7B16 | 0.13% | 1.56s | 88.42 |
@@ 5.3.2 Qualitative Analysis
Examples of Science Question Answering. In Fig. 4, we visualize the examples of EAS-7B12 and LaVIN-7B for both language-only (left) and image-language questions (right). From these examples, we can see that both EAS-7B and LaVIN-7B correctly answer the question. Meanwhile, they all inherit the strong language ability of LLaMA [50], and can explain the answers fluently. Notably, for the visual question (right), EAS-7B has better and more detailed explanations for the answer, e.g., “The colony’s name comes from the fact that it was founded by people from Hampshire, a county in England”. In contrast, the response of LaVIN is logically incorrect even through its answer is right. Overall, these results confirm again the merits of EAS for the efficient adaption of MLLMs.
6 Limitation
EAS focuses on computation-efficient tuning of MLLMs, thus it has no obvious advantages over PETL methods in parameter efficiency. In this case, EAS is superior in the overall target of PCETL.
@@ 7 Conclusion
In this paper, we propose a novel method for parameter and computation efficient tuning of Multi-modal Large Language Models (MLLMs), named by Effective Attention Skipping. Concretely, we first reveal that not all MHAs are necessary for the efficient adaption of MLLMs, based on which EAS adopts a granular redundancy evaluation scheme. Meanwhile, to avoid the additional computation caused by the adapter-based skip connections, EAS is also equipped with a novel Propagation-of-Information Adapter (PIA), which can not only keep parameter efficiency but also can be re-parameterized into the model without extra latency. To validate EAS, we apply it to a recent MLLM called LaVIN, and conduct extensive experiments on the multi-modal benchmark ScienceQA. The experimental results show that EAS can achieve better performance than LaVIN while speeding up inference by up to 2.18 times.
Appendix A Appendix
A.1 Influence of hidden dimension
@@ A.3 The Detailed Results
| LaVIN-7B [38] | 3.8M | 32.32G | 3.70s | 89.25 | 94.94 | 85.24 | 88.51 | 87.46 | 88.08 | 90.16 | 88.07 | 89.41 |
| DAS-7B2 [53] | 4.2M | 32.44G | 3.44s | 88.68 | 94.94 | 86.45 | 88.03 | 86.81 | 88.92 | 90.20 | 88.00 | 89.41 |
| DAS-7B4 [53] | 4.6M | 31.67G | 3.23s | 88.99 | 94.60 | 85.09 | 87.88 | 86.51 | 88.36 | 89.72 | 88.13 | 89.15 |
| DAS-7B6 [53] | 5.0M | 30.91G | 3.06s | 87.30 | 93.36 | 82.36 | 86.12 | 85.97 | 85.71 | 88.18 | 85.70 | 87.29 |
| DAS-7B6 [53] | 5.4M | 30.05G | 2.94s | 86.50 | 89.31 | 84.27 | 85.63 | 83.74 | 86.55 | 87.41 | 84.90 | 86.51 |
| EAS-7B2(Ours) | 4.5M | 32.82G | 2.13s | 88.81 | 95.16 | 85.36 | 87.29 | 86.32 | 88.99 | 90.35 | 87.28 | 89.25 |
| EAS-7B4(Ours) | 5.1M | 32.50G | 2.08s | 88.81 | 94.83 | 86.36 | 87.68 | 86.81 | 89.06 | 90.42 | 87.67 | 89.44 |
| EAS-7B6(Ours) | 5.7M | 32.05G | 1.99s | 88.99 | 94.15 | 85.36 | 87.88 | 87.46 | 88.92 | 90.68 | 86.68 | 89.35 |
| EAS-7B8(Ours) | 6.4M | 31.63G | 1.92s | 90.05 | 94.04 | 85.82 | 89.25 | 87.56 | 88.78 | 90.68 | 88.20 | 89.79 |
| EAS-7B12(Ours) | 7.7M | 31.10G | 1.69s | 89.52 | 95.61 | 86.36 | 88.71 | 88.25 | 88.85 | 91.26 | 87.67 | 89.98 |
@@
A.4 The detailed skipping results
We first visualize the accelerated MLLM searched by DAS [53] and our EAS. From Fig. 6, we can observe that the layers skipped by both DAS and EAS tend to be deep layers. While the layers that are considered to have redundancy in EAS are deeper than those in DAS. Furthermore, we can find out that the skipped modules are concentrated behind the middle of MLLM, and gradually expand forward as the number of skipped modules increases.